# Systematic Study of Amorphous ABC Heterostructures at the Atomic Scale as a Nonlinear Optical Metamaterial

### v1

@@ IV Conclusion
We demonstrated an enhancement of the effective bulk second-order susceptibility in ABC type heterostructures by optimizing the thickness of the ABC period, achieving an optimal value of $t_{\text{ABC}}=1.5$ nm and a dominant tensor component of $\chi_{zzz}=2.0\pm 0.2$ pm/V, based on the second-order nonlinear response at the interfaces. The observed non-linearity enhancement stems from the higher density of interfaces, symmetry breaking in the ABC stacks, and dielectric contrast between the materials. The nonlinearity could be further enhanced if the layers remain distinct and non-intermixed at sub-0.5 nm thicknesses, for example, by minimizing surface and interface roughness through deposition on ultra-smooth substrates. Experimental results indicate that the optimal strategy for maximizing second-order susceptibility is to incorporate the greatest number of ABC layers per unit thickness, while ensuring physical separation to maintain structural integrity. However, the layer thickness is ultimately constrained by the limit of a single atomic layer. Therefore, future efforts should prioritize identifying a more suitable material composition, which remains challenging and elusive. Additionally, ABC nanolaminates have potential for near-ultraviolet applications, leveraging their low absorption edge when constituent materials are carefully selected. Polarity flipping, combined with form birefringence, enables phase matching in bulk form-birefringent metamaterials formBirometamaterial . We envision that ABC type heterostructures will open new avenues for nonlinear optics by integration into nanophotonic waveguides to efficiently access the dominant tensor component $\chi_{zzz}$ ABC-waveguide , metasurfaces and other photonic platforms ABC-overcoating ; ABC-metasrufaceplasmonic .
Funding
@@
This research received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the International Research Training Group (IRTG) 2675 “Meta-ACTIVE” (project number 437527638) and the Collaborative Research Center (CRC) 1375 “NOA.” Additional support was provided by the Fraunhofer Internal Programs under Grant No. SME 431 40-04871. We also gratefully acknowledge funding from the Federal Ministry of Research, Technology and Space under Project ID 13N16897 (Gradient). Authors also thank the European Regional Development Fund and the State of Brandenburg for the Themis Z microscope (part of Potsdam Imaging and Spectral Analysis (PISA) facility).
Disclosures
The authors declare no conflicts of interest.
Data Availability
@@
Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
Appendix A Appendix
The following method On-determination-SHG was used to experimentally obtain the temporal walk-off time for our laser and 1 mm thick fused silica substrates. The femtosecond laser with p-polarization of the fundamental at a power of 1 W was used to measure the angle-dependent p-polarized SH signal from the front and back surfaces with known nonlinear susceptibility $\chi_{\textup{s}}^{\textup{glass}}$ on each side. Then, the equation is as follows:
$=-j\frac{\omega}{2cN_{\textup{glass}}\cos(\Theta_{\textup{glass}})}t^{2}_{\textup{air,glass}}T_{\textup{glass,air}}E^{2}_{\omega,\textup{in}}$
$\ \ \ \ \ \left(\chi_{\textup{s},xxz}^{\textup{glass}}\sin(2\theta_{\textup{glass}})\cos(\Theta_{\textup{glass}})+\chi_{\textup{s},zxx}^{\textup{glass}}\sin(\Theta_{\textup{glass}})\cos^{2}(\theta_{\textup{glass}})+\chi_{\textup{s},zzz}^{\textup{glass}}\sin^{2}(\theta_{\textup{glass}})\sin(\Theta_{\textup{glass}})\right)$
$\ \ \ \ \Bigg[\exp\left(-j\frac{2\omega N_{\textup{glass}}\cos(\Theta_{\textup{glass}})L_{\textup{glass}}}{c}\right)-\exp\left(-j\frac{2\omega n_{\textup{glass}}\cos(\theta_{\textup{glass}})L_{\textup{glass}}}{c}\right)\Bigg].$
After substituting into the Eq. 6, we fit the parameters $K_{2}$, $L_{\textup{glass}}$ and $t_{\text{walk-off}}$. The fitting is restricted to the angle of incidence of 60° to limit the effect of multiple reflections in the sample. The obtained value for the temporal walk-off time is $t_{\text{walk-off}}=94$ fs, see Fig. 13.
Figure 13: The result of the fitting for the temporal walk-off time $t_{\text{walk-off}}$ for 1 mm fused silica substrate.
Other parameters for the evaluation of the second-order susceptibility are shown in the Table 5.
Table 5: Additional fitted parameters of SiO2/TiO2/Al2O3 heterostructures for the evaluation of the second-order susceptibility
Sample ID Calibration constant $K_{2}$ (-) $\times 10^{3}$ $L_{\text{glass}}$ ($m$)
S1 26 1.039
S2 88 1.010
S3 92 1.019
S4 83 0.999
S5 60 0.985
S6 95 1.043
S7 104 1.048
S8 101 1.027
S9 100 0.985
S10 70 1.025
S11 103 1.017
S12 88 0.999
S13 108 1.026
S14 90 1.017

### v4

@@ 4 Conclusion
We demonstrated an enhancement of the effective bulk second-order susceptibility in ABC-type heterostructures by optimizing the thickness of the ABC period, achieving a dominant tensor component of $\chi_{zzz}=2.0\pm 0.2$ pm/V at $t_{\text{ABC}}=1.5$ nm, based on the second-order nonlinear response at the interfaces. The observed non-linearity enhancement stems from the higher density of interfaces, symmetry breaking in the ABC stacks, and dielectric contrast between the materials. The nonlinearity could be further enhanced if the layers remain distinct and non-intermixed at the range of 0.5 nm, for example, by minimizing surface and interface roughness through deposition on ultra-smooth substrates. Extremely thin and homogeneous layers are needed. The evolution of the roughness inside the ALD multilayer stack is the result of both the replicated substrate roughness and the intrinsic film roughness. The coating material and process conditions can influence the latter. This may have the potential for further improvements, particularly in increasing layer density [34]. Among other techniques, ALD has a high potential to provide such films for future advanced nonlinear photonic application. Experimental results indicate that the optimal strategy to maximize second-order susceptibility is to incorporate the greatest number of ABC layers per unit thickness, while ensuring physical separation to maintain structural integrity. However, the layer thickness is ultimately constrained by the limit of a single atomic layer. Therefore, future efforts should also prioritize the identification of a more suitable material composition, which remains challenging and elusive. Additionally, ABC nanolaminates have potential for near-ultraviolet applications, leveraging their low absorption edge when constituent materials are carefully selected. Polarity flipping, combined with form birefringence, enables phase matching in bulk form-birefringent metamaterials [39]. We envision that ABC-type heterostructures will open new avenues for nonlinear optics through integration into nanophotonic waveguides to efficiently access the dominant tensor component $\chi_{zzz}$ [12], and current efforts focus on integrating the heterostructures into nonlinear metasurfaces and other photonic platforms [16, 30].
[…]
@@
Funding This research received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the International Research Training Group (IRTG) 2675 “Meta-ACTIVE” (project number 437527638) and the Collaborative Research Center (CRC) 1375, (project number 398816777) "NOA". Additional support was provided by the Fraunhofer Internal Programs under Grant No. SME 431 40-04871. We also gratefully acknowledge funding from the Federal Ministry of Research, Technology, and Space under Grant No. 13N16897 (GRADIENT) and 13F1000A (PriFusio). The authors also thank the European Regional Development Fund and the State of Brandenburg for the Themis Z microscope (part of the Potsdam Imaging and Spectral Analysis (PISA) facility).
[…]
Disclosures The authors declare no conflict of interest.
[…]
@@
Data Availability Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
[…]
Supplemental document See Supplement 1 for additional details of the nonlinear characterization.